OpenStudy (anonymous):

Sherlock wants to rent a textbook. He has to pay a fixed base cost plus a daily rate for renting the textbook. The table shows the amount of money, y, in dollars, that Sherlock has to pay for renting the textbook for x days number of days(x) rent(dollars)(y) 0 6 1 13 2 20 3 27 4 34

OpenStudy (anonymous):

Which equation best shows the relationship between x and y? y=x+13 y=7x+13 y=x+6 y=7x+6

OpenStudy (anonymous):

@DIDITHEDAD can you help? ill medal and fan.

OpenStudy (anonymous):

@tylermcmullen23

OpenStudy (anonymous):

@TylerD

OpenStudy (anonymous):

well, a) 6=13??? nope b)6=13???? nope c)6=6??yep, 13=7? nope, so its incorrect d)6=6?? yep! 13=13, yep! so its correct

OpenStudy (anonymous):

thank you! medal!!! and fan!!